Observe yourself in a day. Find out how many hours you spend in the following 

activities: house chores, answering Self Learning Modules, planting/gardening, 

using social media like Facebook, messenger, tiktok, Instagram, and YouTube, 

listening to music, watching television, and sleeping. Record your data. Construct a 

probability distribution, then compute the mean (µ), variance (σ2), and standard 

deviation (σ) of the probability distribution that you made. 


1
Expert's answer
2021-05-07T08:52:43-0400

The probability Distribution is as follows-



Mean of Probability Distribution-

"E(X)=\sum X.P(X)=4.6596"


Variance of Probability distribution-

"Var=E(X^2)-[E(X)]^2=26.265-(4.6596)^2=26.265-21.71=4.55"


Standard deviation of probability distribution-

"S.D.=\sqrt{Var}=\sqrt{4.55}=2.13"